OpenStudy (anonymous):

Question: 42 of 66: Anthony shines a beam of light on the surface of a solution of glycerol at a certain angle. If the angle of refraction is 38°, what is the angle of incidence? The index of refraction of glycerol is 1.47, while that of air is 1.00. Select one of the options below as your answer: A. 38° B. 45° C. 55° D. 65° E. 70°

OpenStudy (fifciol):

 \[\frac{ \sin \alpha }{ \sin \theta }=\frac{ n_{glycerol} }{ n_{air} }\] so \[\sin \alpha =1,47* \sin 38^{o} \rightarrow \alpha =65^{o}\] Answer D